OpenStudy (anonymous):

PLEASE HELP: Suppose that in 2001, the average cost of a gallon of milk was $2.43. In 2007 the average price was $4.21. Assuming the relationship is linear, find the equation of the line that models the price of milk over time since 1999. Round to the nearest thousandth.

OpenStudy (anonymous):

( 2001, 2.43) ( 2007, 4.21) Can you find slope of these 2 points?

OpenStudy (anonymous):

First change your independent variables (Year) into years since 1999 2001 -1999 = 2 2007 - 1999 = 8 Then you use slope, and then solve for b

OpenStudy (anonymous):

@Chlorophyll Slope: 1.78/6

OpenStudy (anonymous):

&& what do I do next

OpenStudy (anonymous):

y - y' = m ( x - x' ) Suppose you use ( 2001, 2.43) to plug in

OpenStudy (anonymous):

& 2001 , 2.43 being x1 and y1? or no

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

just leaving the y alone and the x, sorry i really dont know

OpenStudy (anonymous):

Yes, after distribute and simplify you got the linear line y = mx + b